jabez177 (jabez177):

Mitchell poured the contents of a completely filled cone into an empty cylinder, and the cylinder became two-thirds full. The cylinder had a radius of 5 cm and a height of 15 cm. What were possible dimensions of the cone? Use 3.14 to approximate pi. A. h = 5 cm; r = 15 cm B. h = 7.5 cm; r = 10 cm C. h = 10 cm; r = 7.5 cm D. h = 15 cm; r = 5 cm

jabez177 (jabez177):

@FortyTheRapper @raffle_snaffle

ILovePuppiesLol (ilovepuppieslol):

hi

OpenStudy (fortytherapper):

Let's first find the volume of the cylinder That's \[\pi*r^2*h\]

ILovePuppiesLol (ilovepuppieslol):

oh this is alot of work we need formulas and stuff

jabez177 (jabez177):

1, 177.5 is the volume of the cylinder.

OpenStudy (fortytherapper):

Right Since it only filled up 2/3rds of the cylinder though, that means \[(1177.5)(2/3) = V _{cone}\]

jabez177 (jabez177):

Thought so...

OpenStudy (raffle_snaffle):

@jabez177 hey it's good to double check with someone else. Good job.

jabez177 (jabez177):

785 is the volume of the cone.

OpenStudy (fortytherapper):

Right, so that means \[\frac{ 1 }{ 3 }*\pi*r^2*h = 785\] So now it's a matter of getting R and H by itself, like they were x's in an equation

OpenStudy (fortytherapper):

\[\frac{ 1 }{ 3 }*3.14*r^2*h = 785\] There we go, so numbers on one side, variables on the other

jabez177 (jabez177):

alright

jabez177 (jabez177):

where do we get the radius and stuff from tho?

OpenStudy (fortytherapper):

The answer choices We isolate to get that r^2 + h = value We then use those answer choices to plug in and see if it satisfies that value

jabez177 (jabez177):

B.

OpenStudy (fortytherapper):

Perfecto

jabez177 (jabez177):

Thanks! :)